OpenStudy (anonymous):

4(7 - 3x) = 7(4 - 2x)

OpenStudy (saifoo.khan):

28 - 12x = 28 - 14x

OpenStudy (saifoo.khan):

x = 0

OpenStudy (aroub):

28-12x=28-14x -12x+14x=-28+28 2x=0 x=0